# Impact of the pre-equilibrium phase for the determination of nuclear geometry in high-energy isobar collisions

### v2

@@ IV Conclusion
We investigated the effects of a prethermalization phase in isobar collision systems (namely ${}^{96}_{44}$u +${}^{96}_{44}$u and ${}^{96}_{40}$r +${}^{96}_{40}$r ), using free-streaming as an extreme case. The main effects anticipated were observed: the smearing out of small-scale features and reduction of spatial anisotropies (which get transferred to momentum distribution anisotropy). Despite the absolute magnitude of these effects being large, it is not the same when doing ratios for a given $\tau_{FS}$. In this case, for ratios of ellipticity and rms radii, differences are less than 1% and unlikely to be relevant when comparing to data. Ratios of higher order eccentricities, however, show differences that can be significant when comparing with precision measurements such as the ones recorded for isobar collisions. Within our statistics, not much can be learned from the two-plane correlator for elliptic and triangular flows but interestingly the normalized symmetric cumulant seems sensitive to the duration of free-streaming. These studies of eccentricities yield information on the parameters $\beta_{2}$, $\beta_{3}$ and $a$.
Furthermore, we studied the effects that nuclear geometry could have on the correlation between anisotropic flow and mean momentum to get information on $\gamma$. In Giacalone et al. (2021c), the estimator $\rho(\epsilon_{2}^{2},E/S)$ was shown to better reproduce the hydrodynamic correlator $\rho(v_{2}^{2},[p_{T}])$. Here for the first time, it is used to study isobar collisions. A several percent difference between the cases n=1 and 2 (with and without triaxiality) is indeed observed for $\rho_{2}$. In addition, and this is the main result as far as free-streaming effects are concerned, the duration of the prethermalization phase, regulated by $\tau_{FS}$, has a noticeable and diverse effect on these ratios of correlators depending on the description of the nuclear geometry: for $\rho_{2}$, it is able to enhance the separation between results with and without triaxiality effects (cases 1 and 2); for $\rho_{3}$ it generates a non-zero correlation between $\varepsilon_{3}$ and $E/S$ (proxies for $v_{3}$ and the mean transverse momentum, respectively) for the same cases. Another distinct effect for $\rho_{2}$ appears for case 3 (where in addition to the quadrupole deformation of cases 1 and 2, an octupole deformation is introduced), a double sign change for the largest free-streaming times ($1.35\lesssim\tau_{FS}\,({\rm fm/c})\lesssim 1.49$) appears, which can be contrasted to no or perhaps one sign change for initial condition and free-streaming time equal to $1$ fm/c. If verified experimentally, these features provide another tests of the predictive power of state-of-the-art hybrid hydrodynamic models Giannini et al. (2023).

### v4

@@ IV Conclusion
We investigated the effects of a pre-thermalization stage in isobar collision systems (namely ${}^{96}_{44}$u +${}^{96}_{44}$u and ${}^{96}_{40}$r +${}^{96}_{40}$r ), using free-streaming as an extreme case. The main effects anticipated were observed: the smearing out of small-scale features and reduction of spatial anisotropies (which get transferred to momentum distribution anisotropy). Despite the absolute magnitude of these effects being large, it is not the same when doing ratios for a given $\tau_{FS}$. In this case, for ratios of ellipticity and rms radii, differences are less than 1% and unlikely to be relevant when comparing to data. Ratios of higher order eccentricities, however, show differences that can be significant when comparing with precision measurements such as the ones recorded for isobar collisions. Within our statistics, not much can be learned from the two-plane correlator for elliptic and triangular flows but interestingly the normalized symmetric cumulant $\varepsilon NSC(2,3)$ seems sensitive to the duration of free-streaming. These studies of eccentricities yield information on the parameters $\beta_{2}$, $\beta_{3}$ and $a$.
Furthermore, we studied the effects that nuclear geometry could have on the correlation between anisotropic flow and mean momentum to get information on $\gamma$. In [52], the estimator $\rho(\varepsilon_{2}^{2},E/S)$ was shown to better reproduce the hydrodynamic correlator $\rho(v_{2}^{2},[p_{T}])$. Here for the first time, it is used to study isobar collisions. A several percent difference between the configurations $i=1$ and 2 (with and without triaxiality) is indeed observed for $\rho_{2}$. In addition, and this is the main result as far as free-streaming effects are concerned, the duration of the pre-thermalization stage, regulated by $\tau_{FS}$, has a noticeable and diverse effect on these ratios of correlators depending on the description of the nuclear geometry: for $\rho_{2}$, it is able to enhance the separation between results with and without triaxiality effects (configurations 1 and 2); for $\rho_{3}$ it generates a non-zero correlation between $\varepsilon_{3}$ and $E/S$ (proxies for $v_{3}$ and the mean transverse momentum, respectively) for the same configurations. Another distinct effect for $\rho_{2}$ appears for configuration 3 (where in addition to the quadrupole deformation of configurations 1 and 2, an octupole deformation is introduced), a double sign change for the largest free-streaming times ($1.35\lesssim\tau_{FS}\,({\rm fm/c})\lesssim 1.49$) appears, which can be contrasted to no or perhaps one sign change for initial condition and free-streaming time equal to $1$ fm/c. If verified experimentally, these features provide another tests of the predictive power of state-of-the-art hybrid hydrodynamic models [57].